OpenStudy (anonymous):

Consider the following pair of equations: y = x + 4 y = −2x − 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

pooja195 (pooja195):

it already tells u the value of y so plug it in opposite equations :) E1 -----> E2

pooja195 (pooja195):

\[\huge~ x + 4= −2x − 2 \] Now you have to combine like terms

pooja195 (pooja195):

x+2x=? -2-4=?

OpenStudy (anonymous):

3x and -6

pooja195 (pooja195):

good \[\huge3x=-6\] Whats -6/3

OpenStudy (anonymous):

-2

pooja195 (pooja195):

Good so we have x=-2 now just plug in X value into E1 :) \[y = (-2)+ 4\]

OpenStudy (anonymous):

wait im confused

OpenStudy (anonymous):

how did you get that equation?

OpenStudy (anonymous):

OHH

OpenStudy (anonymous):

i understand now :)

pooja195 (pooja195):

good :) so whats the answer?

OpenStudy (anonymous):

y=2

pooja195 (pooja195):

good so we have x=-2 y=2 :)

OpenStudy (anonymous):

Thank you so much!

pooja195 (pooja195):

No problem :)